Triangle RST is isosceles with base RS. If RT = 5x-4, ST = 3x+4, and RS = 2x+9, find the lengths of the sides of the triangle.
Answer by mananth(16949)

In triangle RST RS = base = 2x+9
RT=ST=congruent sides
RT =5x-4
ST=3x+4
3x+4=5x-4
2x=8
x=4
RS= 2*4+9=17
RT=16
ST=16